Calculate K for the reaction :


2A + B <----> C + D


If at equilibrium, A = 0.4 mols/liter, mols/liter, B = 0.5 mols per liter. C = 0.3 mols per liter, and D = 0.8 mols per liter.

1
Expert's answer
2022-03-24T12:13:05-0400

Solution:

The chemical reaction:

2A + B ⇌ C + D

The equilibrium constant expression for the reaction system:



[A] = 0.4 mol/L

[B] = 0.5 mol/L

[C] = 0.3 mol/L

[D] = 0.8 mol/L

Therefore,

K = (0.3 × 0.8) / (0.5 × 0.42) = (0.24) / (0.08) = 3.0

K = 3.0 L/mol


Answer: The equilibrium constant (K) for the reaction is 3.0 liter/mol